The length of a rectangle is 2 less than 3 times the width. If the area of the rectangle is 96 square inches, what is the length of the rectangle?
A)5 1/3 inches
B)6 inches
C)14 inches
D)16 inches

Answer by jorel1380(3719)

LxW=A
(3w-2)w=96
3w2-2w-96=0
(3w+16)(w-6)=0
w=-16/3 or 6
Throwing out the negative answer, we get the width to be 6 inches. Then our length is 16..